7. Let A, B, and C be three events. Write out expressions for the following events in terms of A, B, and C, using set notation. (a) A occurs but neither B nor C occurs. (b) A and B occur, but not C (c) A or B occurs, but not C. (d) None of A, B, and C occurs. (e) A, B, or C occurs (n) Exactly one of A, B, and C occurs (g) Exactly two of A, B, and C occur. (h) Two or more of A, B, and C occur. (I) Less than two of A, B, and C occurs. O Either A occurs and not B, or B occurs and not A.
Answer #1

A, B and C are three events then following are the required expressions-

a.)(A \cap {B}'\cap {C}')=A-(A\cap B)- (A\cap C) +(A\cap B \cap C)\\ b.)(A\cap B \cap {C}')= (A\cap B)-(A\cap B \cap C) \\ c.)(A\cup B )\cap {C}' =(A\cup B \cup C) -C\\ d.)({A}'\cap {B}' \cap {C}')=0\\ e.)(A\cup B \cup C)=A+B+C-(A\cap B)-(B\cap C)-(A\cap B)+(A\cap B\cap C)\\ f.) (A\cap {B}' \cap {C}')+(B\cap {C}' \cap {A}')+(C\cap {B}' \cap {A}')=[A-(A\cap B)- (A\cap C) +(A\cap B\cap C)]+[B-(B\cap A)- (B\cap C) +(B\cap A\cap C)]+[C-(C\cap B)- (C\cap A) +(C\cap B \cap A)]\\ g.)[(A\cap B \cap{C}'+(A\cap C \cap{B}'+(C\cap B \cap{A}']=[(A\cap B)-(A\cap B \cap C)+(A\cap B)-(A\cap C \cap B)+(C\cap B)-(C\cap B \cap A)]\\